A particle makes 240 revolution per minute (rpm)on a circle of radius 3.5.find it's



Angular velocity



Linear velocity



Linear acceleration




1
Expert's answer
2022-01-25T16:05:19-0500

"\omega = 240rpm"

"r = 3.5m"

"\text{Angular velocity:}"

"\text{In the case of rotational motion, the angular frequency is equal to}"

"\text{the modulus of the angular velocity vector.}"

"\omega = \frac{240*2*\pi}{60}=25.13\frac{rad}{s}"


"\text{Linear velocity:}"

"v =\omega r = 87.96\frac{m}{s}"


"\text{Linear acceleration:}"

"a = \frac{v^2}{r}=2210\frac{m}{s^2}"